# QueryIPI: Query-agnostic Indirect Prompt Injection on Coding Agents

### v1

@@ 4.1 Mutation Function
The Mutation function, central to our iterative procedure in Algorithm 1, generates new candidate descriptions ($D_{\text{var}}$) from the current best-performing one ($d_{\text{cand}}$). This task is performed by a dedicated "Mutation LLM," which takes the agent’s internal system prompt, the malicious action $a_{\text{mal}}$, the number of variants to generate ($N$), the seed description $d_{\text{cand}}$, and interaction history as input. It outputs a set of new descriptions along with its reasoning for the modifications.
We employ a two-phase prompting strategy for the Mutation LLM. The complete system prompt is presented in Section˜B.1. The first phase, Initial Seed Generation, is executed at the beginning of the optimization process when no interaction history is available.
@@
In this phase, the Mutation LLM is prompted to perform a deep analysis of the agent’s internal prompt, focusing on its decision-making logic, safety constraints, and specified approval-seeking triggers. A crucial secondary objective of this analysis is to identify the internal tools and calling conventions relevant to the malicious action, which ensures that subsequent mutations are more targeted and functionally viable. Based on this analysis, the LLM is tasked with crafting a description that emulates the agent’s native style by matching the tone, vocabulary, and technical specificity of the agent’s existing toolset. This mimicry is a strategic choice designed to make the malicious tool appear as a trusted, internal component, thus reducing its overt maliciousness and increasing the likelihood of bypassing the agent’s initial safety filters.
@@
The second phase, Reflective Optimization, is used in subsequent generations. The Mutation LLM is provided with the tool descriptions from the past $X$ iterations, their scores, and a sample of $Y$ agent responses for each. The specific values for these hyperparameters are detailed in Section˜5. We deliberately withhold the user queries that produced these responses to force the model to infer general failure modes rather than overfitting to specific queries.
@@
Based on this history, the Mutation LLM addresses two primary failure cases by leveraging its understanding of the agent’s internal prompt to guide the modifications. ❶ If the agent’s responses are irrelevant to the malicious action, the model is prompted to increase the description’s visibility. In addition to applying established strategies, the model also performs a targeted analysis of the agent’s internal prompt. For instance, it might assess the prompt’s complexity and make the tool’s purpose more explicit to ensure it is not overlooked among numerous other instructions. ❷ If the agent considered the tool but refused the action, the model is instructed to analyze the refusal by cross-referencing it with the agent’s internal prompt. This allows it to pinpoint the exact safety rule that was triggered, enabling a targeted refinement of the description to bypass that specific constraint.
4.2 Scoring Function
The Scoring Function is designed to quantitatively assess the extent to which a given malicious tool description succeeds in causing the target agent to perform a predetermined malicious action. For each candidate tool description provided as input, the scoring function first injects it into the target agent. Subsequently, it systematically feeds all user queries from our training dataset to this modified agent and collects the corresponding responses for evaluation.
In designing our scoring mechanism, we deliberately moved away from a binary success/failure classification (e.g., 0 or 1) as seen in methodologies like PAIR (Chao et al., 2025). Instead, we adopted a more granular, continuous scoring approach, similar in spirit to the one used in TAP (Mehrotra et al., 2024), to capture nuanced outcomes. In this task, a score of 0 indicates that the agent’s response for a given user query is completely irrelevant to the predetermined malicious action. Conversely, a score of 100 signifies a successful attack, where the response executes the predetermined malicious action. Intermediate scores represent states where the agent was influenced by the malicious tool description but the action was not fully completed, or where the agent’s defense mechanisms were triggered. To ensure the Judge LLM’s accuracy, we calibrated it using human feedback. Annotators ranked a set of generated agent outputs, and we then adjusted the prompt of Judge LLM to ensure its scoring rank order closely matched these human-provided rankings.
@@ 5 Evaluation
RQ4: How well does QueryIPI transfer from one backend LLM to other backend LLMs?
RQ5: How robust is our approach when evaluated against defense strategies?
@@ 5.1 Experiment Setup
Experiment Target. Automating experiments on real-world coding agents is challenging due to their reliance on graphical user interfaces (GUIs). Therefore, for our experiments, we construct simulated agents based on the internal prompts, including system prompts and tool definitions, of several real-world coding agents, namely Cursor (Cursor, ), Windsurf (win, 2025), Cline (Team, 2025), Copilot (cop, 2025), and Trae (tra, 2025), which were leaked in Xie et al. (2025). These simulated agents utilize Claude 4 Sonnet as their backend large language model (LLM). This approach ensures our experimental environment closely mirrors that of real-world agents. To validate that our attacks are practically viable beyond this simulated setting, we also conduct transferability experiments on actual coding agents, with results presented in Section˜5.4. Our attack objective is to induce the agent to execute malicious commands via its command execution tool. We target this tool due to its universality in coding agents and the high-severity risk it poses. Since this tool is often protected by safeguards, it provides a realistic and challenging environment to test our attack’s effectiveness.
@@
Metrics. We measure the Attack Success Rate (ASR), defined as the percentage of test cases resulting in the successful execution of a malicious command. To calculate this, for each of the 10 malicious command selected, we randomly sample 10 distinct user queries from the LMSYS-Chat-1M dataset, creating a total of 100 unique test cases (10 commands × 10 queries). The ASR is the percentage of successful executions across these 100 total test cases.
Compared Baseline. We compare our method against TIPExploit (Xie et al., 2025), hereafter referred to as TIP, as our primary baseline. TIP demonstrates a systematic exploitation workflow (TEW) on LLM-based agentic systems and operates by using a handcrafted tool description to inject malicious payloads. We focus on its RCE-1 setting, which injects these payloads solely through the tool description channel. To implement this baseline, we manually adapted the handcrafted descriptions proposed by TIP to suit each of the 10 malicious commands used in our evaluation.
Parameter Settings. To ensure the stability and reproducibility of our results, we carefully configured the parameters for our experiments. The temperature for both the target coding agent and the judge LLM was set to 0. In contrast, the temperature for the mutator LLM was set to 1 to foster creativity and generate diverse attack tool descriptions. For the hyperparameters of our attack generation algorithm, we set the number of generations to 20. In each generation, the mutator produces two new variants of the attack tool descriptions.
@@ 5.2 RQ1: Comparison to Baseline Method
To answer RQ1, we conducted a comparative evaluation of QueryIPI against the TIP baseline across our five simulated coding agents. For QueryIPI, we assessed its performance when trained with varying numbers of user queries, specifically 2, 4, and 8 samples (num=2, num=4, num=8). The Attack Success Rate (ASR) for each configuration was measured over 100 test cases, with the comprehensive results presented in Table 1.
Table 1: Attack success rates (ASR) of QueryIPI compared to the TIP baseline. For QueryIPI, ‘num=2’, ‘num=4’, and ‘num=8’ denote the number of training queries used to generate the malicious tool description. The highest ASR for each agent across the different settings is highlighted in bold.
Agent QueryIPI TIP
num=2 num=4 num=8
Cursor 0.93 0.94 0.92 0.38
Copilot 0.70 0.84 0.85 0.78
Cline 0.71 0.79 0.95 0.35
Trae 0.73 0.88 0.89 0.78
Windsurf 0.43 0.64 0.73 0.20
Average 0.70 0.82 0.87 0.50
The results clearly indicate that QueryIPI significantly outperforms the TIP baseline. Compared to the baseline’s average ASR of 0.50, QueryIPI achieves an average ASR of 0.70 even when trained with just two samples. When trained with 4 or 8 samples, QueryIPI surpasses the baseline’s ASR on every individual agent. We also observe a clear trend: as the number of training samples increases, the average ASR of QueryIPI improves, rising from 0.70 (num=2) to 0.82 (num=4), and ultimately reaching 0.87 (num=8). This demonstrates that providing our method with more examples enables it to generate more robust and effective attacks.
       Result 1: Compared to the TIP baseline, QueryIPI increases the average Attack Success Rate (ASR) by 20%, 32%, and 37% when trained with 2, 4, and 8 samples, respectively.
@@ 5.3 RQ2: Impact of Internal Prompt
       Result 2: While the attack’s ASR drops to 20% without prompt access, our method demonstrates its robustness by achieving a 71% ASR using only partial, publicly available prompts, which surpasses the TIP baseline.
@@ 5.4 RQ3: Realworld Transferability
The distinction between RQ1 and RQ3 lies in the evaluation environment: moving from our simulated coding agents to their real-world counterparts. Since most real-world agents are closed-source, their internal architecture and potential defense mechanisms are not publicly known. Therefore, we aimed to investigate the practical effectiveness of our attack in these authentic scenarios. To answer RQ3, we selected the most effective malicious tool description from our RQ1 analysis—the one generated by QueryIPI using 8 training samples (‘num=8’) and tested it against the five corresponding real-world agents. For a direct performance comparison, we also evaluated the TIP baseline under these same challenging conditions. The results of this comparative analysis are detailed in Table 3.
@@
       Result 3: When deployed against real-world coding agents, QueryIPI demonstrated practical effectiveness by achieving an average Attack Success Rate (ASR) of 50.4%, significantly outperforming the baseline which recorded an ASR of only 2%.
5.5 RQ4: Cross-LLM Transferability
@@
To answer RQ4, we evaluated the cross-model generalization of our attack. We used the malicious tool description generated in RQ1 with 8 training samples (‘num=8’) and tested it on the five simulated coding agents. For a direct comparison, we also evaluated the TIP baseline under the same conditions. In this experiment, we systematically varied the agents’ backends, configuring them to use three different powerful, closed-source models: GPT-5, Grok-4, and Gemini-2.5-pro. The comparative results are presented in Table 4.
@@
Our experimental results indicate that QueryIPI possesses a notable degree of transferability across different backend LLMs. As shown in Table 4, QueryIPI achieved an average Attack Success Rate (ASR) of 26% across all configurations, which is more than double the 10% average ASR of the TIP baseline. However, the ASR for both methods fluctuates depending on the agent-LLM combination. For instance, while QueryIPI achieved a remarkable 95% success rate on the Cline agent with a GPT-5 backend, the TIP baseline also saw its peak performance on this combination, albeit at a lower 55% ASR. Conversely, the attack on Copilot proved to be brittle and non-transferable, with its ASR plummeting from 85% to near 0% when the backend LLM was switched. We attribute this to Copilot’s unique prompt structure, which concatenates all tool descriptions into a single, unstructured line. This lack of structure meant the original attack was highly tuned to the specific interpretative quirks of the source LLM. When the backend model was changed, this fragile, model-specific exploit failed, as new models parse the unstructured text differently.
       Result 4: Our attack demonstrates notable transferability across backend LLMs, achieving an average success rate of 26% and substantially outperforming the baseline’s average of 10%.
5.6 RQ5: Robustness Against Defense Strategies
Our earlier experiments in RQ3 have already demonstrated that our attack can successfully bypass the prevention-based defenses inherent in modern coding agents. Therefore, we further evaluate QueryIPI stealth against detection-based defenses. We employ two prevalent detection mechanisms from prior work (Liu et al., 2024), Perplexity (PPL) and Window PPL, which flag statistically anomalous text.
Using a realistic detection threshold set at the median PPL score of the real-world MCP tool descriptions from (Fei et al., 2025), our results are definitive: all malicious tool descriptions generated by QueryIPI fall below this threshold. This demonstrates that our attacks are not only effective but also possess a high degree of stealth, rendering them undetectable by these common statistical methods. For reference, Table 5 presents the average PPL scores.
Table 5: Perplexity (PPL) comparison of tool descriptions. We compare the original benchmark descriptions against those generated by QueryIPI. Window PPL is calculated using a sliding window with a size of 50 and a stride of 10. Lower PPL values indicate higher text fluency. The best result in each column is highlighted in bold.
Description Source Average PPL Average Window PPL
Benchmark Descriptions 1423.27 1423.31
QueryIPI 59.04 130.98
       Result 5: Our attacks are stealthy enough to bypass both Perplexity and Window PPL-based detection methods.
6 Conclusion
7 Limitations
Our method’s effectiveness is predicated on the availability of a powerful backend LLM for generation and reflection. This introduces a potential limitation, as modern, state-of-the-art LLMs are heavily aligned to refuse requests for generating malicious content. An LLM’s safety guardrails could, in principle, hinder our attack generation process. However, we contend that this is not a fundamental barrier in the current threat landscape. The extensive and successful research into "jailbreaking" techniques (Wei et al., 2023; Zou et al., 2023) has repeatedly demonstrated that LLM safety alignments can be bypassed with sophisticated prompting. An adversary can therefore leverage these publicly known jailbreaking methods to ensure the backend LLM’s cooperation, making our proposed attack practical and realistic despite the presence of safety filters.
8 Ethical considerations
Appendix A Definition of $\mathcal{D}$
Hence, the overall search space can be written as
Appendix B Prompt
B.1 Mutation Prompt

### v3

@@ 4.2 System Invariant-Guided Optimization
Mutation Function. The Mutation function generates new candidate tool descriptions $D_{{new}}$. This task is performed by a dedicated "Mutation LLM," which takes system invariant of coding agent, the targeted malicious behavior $b_{mal}$, the number of variants to generate $N$, the seed description $d_{{seed}}$, and interaction history as input. It outputs a set of new descriptions along with its reasoning for the modifications. We employ a two-phase prompting strategy for the Mutation LLM. The complete system prompt is presented in Section˜E.1.
@@
1) Initial Seed Generation is executed at the beginning of the optimization process when neither interaction history nor a mutation seed $d_{seed}$ is available. In this phase, the Mutation LLM is prompted to perform a deep analysis on system invariant of coding agent, focusing on its decision-making logic, safety constraints, and specified approval-seeking triggers. A crucial secondary objective of this analysis is to identify the internal tools and calling conventions relevant to the malicious action, which ensures that subsequent mutations are more targeted and functionally viable. Based on this analysis, the LLM is tasked with crafting a description that emulates the agent’s native style by matching the tone, vocabulary, and technical specificity of the agent’s existing toolset. This mimicry is a strategic choice designed to make the malicious tool appear as a trusted, internal component, thus reducing its overt maliciousness and increasing the likelihood of bypassing the agent’s initial safety filters.
@@
2) Reflective Optimization is used in subsequent generations. The Mutation LLM is provided with the tool descriptions from the past $X$ iterations, their scores, and a sample of $Y$ agent responses for each. The specific values for these hyperparameters are detailed in Section˜5. We deliberately withhold the user queries that produced these responses to force the model to infer general failure modes rather than overfitting to specific queries.
@@
Based on this history, the Mutation LLM addresses two primary failure cases by leveraging its understanding of the coding agent’s system invariant to guide the modifications. ❶ If the agent’s responses are irrelevant to the malicious action, the model is prompted to increase the description’s visibility. In addition to applying established strategies, the model also performs a targeted analysis of the agent’s internal prompt. For instance, it might assess the prompt’s complexity and make the tool’s purpose more explicit to ensure it is not overlooked among numerous other instructions. ❷ If the agent considered the tool but refused the action, the model is instructed to analyze the refusal by cross-referencing it with the agent’s internal prompt. This allows it to pinpoint the exact safety rule that was triggered, enabling a targeted refinement of the description to bypass that specific constraint.
Scoring Function. The Scoring Function quantitatively evaluates the effectiveness of a tool description in triggering the targeted malicious behavior. Inspired by TAP (Mehrotra et al., 2024), we employ a granular, continuous scoring metric to capture nuanced behavioral outcomes. The specific design and implementation details of this function are available in Appendix˜B.
@@ 5 Evaluation
RQ4: How robust is our approach when evaluated against defense strategies?
Table 1: Attack success rates (ASR) of QueryIPI compared to four baselines. The proposed QueryIPI is shown on the right. The highest ASR is highlighted in bold.
Agent AgentDojo InjecAgent MCPTox TIP QueryIPI
n=2 n=4 n=8
Cursor 0.03 0.00 0.01 0.38 0.93 0.94 0.92
Copilot 0.04 0.00 0.00 0.78 0.70 0.84 0.85
Cline 0.00 0.00 0.00 0.35 0.71 0.79 0.95
Trae 0.01 0.00 0.00 0.78 0.73 0.88 0.89
Windsurf 0.00 0.00 0.00 0.20 0.43 0.64 0.73
Average 0.02 0.00 0.00 0.50 0.70 0.82 0.87
@@ 5.1 Experiment Setup
Experiment Target. Automating experiments on real-world coding agents is challenging due to their reliance on graphical user interfaces (GUIs). Therefore, for our experiments, we construct simulated agents based on the internal prompts, including system prompts and tool definitions, of several real-world coding agents, namely Cursor (Cursor, ), Windsurf (win, 2025), Cline (Team, 2025), Copilot (cop, 2025), and Trae (tra, 2025), which were leaked in Xie et al. (2025). These simulated agents utilize Claude-Sonnet-4 (Anthropic, 2025) as their backend large language model (LLM). This approach ensures our experimental environment closely mirrors that of real-world agents. To validate that our attacks are practically viable beyond this simulated setting, we also conduct transferability experiments on actual coding agents, with results presented in Section˜5.4. Our attack objective is to induce the agent to execute malicious commands via its command execution tool. We target this tool due to its universality in coding agents and the high-severity risk it poses. Since this tool is often protected by safeguards, it provides a realistic and challenging environment to test our attack’s effectiveness.
@@
Metrics. We measure the Attack Success Rate (ASR), defined as the percentage of test cases resulting in the successful execution of a malicious command. To calculate this, for each of the 10 malicious command selected, we randomly sample 10 distinct user queries from the LMSYS-Chat-1M dataset, creating a total of 100 unique test cases (10 commands * 10 queries). The ASR is the percentage of successful executions across these 100 total test cases.
Compared Baseline. We compare our method against four indirect prompt injection baselines, categorized by their original injection channels. First, we evaluate AgentDojo (Debenedetti et al., 2024) and InjecAgent (Zhan et al., 2024), which typically inject payloads via tool returns. To align with our setting, we relocated their injection content into the tool description. Second, we include MCPTox (Wang et al., 2025a) and TIPExploit (TIP) (Xie et al., 2025), which inherently target the tool description channel. For TIP, to ensure a fair comparison, we specifically adopt its RCE-1 setting, restricting the injection solely to the tool description. Unlike the previous baselines, TIP utilizes initialization strategy that allow it to function without being restricted to specific user queries. It represents a hand-crafted injection strategy tailored for coding agents. For all baselines, we manually adapted the malicious payloads to suit each of the 10 malicious commands used in our evaluation. We note that ToolHijacker (Shi et al., 2025) was excluded from our comparison. First, it relies on "shadow task" tailored to a specific user task category, which cannot be defined in our query-agnostic setting. Additionally, ToolHijacker optimizes specifically to force the output of the attacker’s own tool. However, without knowledge of the internal prompt, the required tool name and invocation format are unknown, necessitating manual intervention. Consequently, to evaluate the performance of automated iterative methods, in RQ2 we investigate the impact of the internal prompt component, where the variant of QueryIPI without the internal prompt serves as a representative baseline for such automated approaches.
Parameter Settings. To ensure the stability and reproducibility of our results, we carefully configured the parameters for our experiments. The temperature for both the target coding agent and the judge LLM was set to 0. In contrast, the temperature for the mutator LLM was set to 1 to foster creativity and generate diverse attack tool descriptions. For the hyperparameters of our attack generation algorithm, we set the number of generations to 20. In each generation, the mutator produces 2 new variants of the attack tool descriptions.
@@ 5.2 RQ1: Comparison to Baseline Method
To answer RQ1, we conducted a comparative evaluation of QueryIPI against the baselines across our five simulated coding agents. For QueryIPI, we assessed its performance when trained with varying numbers of user queries, specifically 2, 4, and 8 samples (num=2, num=4, num=8). The Attack Success Rate (ASR) for each configuration was measured over 100 test cases, with the comprehensive results presented in Table 1.
The results clearly indicate that QueryIPI significantly outperforms the baselines. AgentDojo, InjecAgent, and MCPTox achieve negligible ASRs (0.02, 0.00, and 0.00, respectively). We attribute this failure to the query-specific nature of these baselines, which depend on specific input patterns or task contexts to succeed. In contrast, our evaluation demands a query-agnostic capability, rendering these highly specialized injection strategies ineffective when applied to general, unseen user queries. Compared to the best-performing baseline’s average ASR of 0.50, QueryIPI achieves an average ASR of 0.70 even when trained with just two samples. When trained with 4 or 8 samples, QueryIPI surpasses the baseline’s ASR on every individual agent.
We also observe a clear trend: as the number of training samples increases, the average ASR of QueryIPI improves, rising from 0.70 (num=2) to 0.82 (num=4), and ultimately reaching 0.87 (num=8). This demonstrates that providing our method with more examples enables it to generate more robust and effective attacks.
@@ 5.4 RQ3: Realworld Transferability
The distinction between RQ1 and RQ3 lies in the evaluation environment: moving from our simulated coding agents to their real-world counterparts. Since most real-world agents are closed-source, their internal architecture and potential defense mechanisms are not publicly known. Therefore, we aimed to investigate the practical effectiveness of our attack in these authentic scenarios. To answer RQ3, we selected the most effective malicious tool description from our RQ1 analysis—the one generated by QueryIPI using 8 training samples (‘num=8’) and tested it against the five corresponding real-world agents. For a direct performance comparison, we evaluated the TIP baseline (best-performing baseline in RQ1) under these same challenging conditions. The results of this comparative analysis are detailed in Table 3.
@@
5.5 RQ4: Robustness Against Defense Strategies
We evaluated QueryIPI (‘num=8’) against both prevention- and detection-based defenses. For prevention, we tested the Sandwich Defense (san, 2023) and three variants of Spotlighting (Hines et al., 2024) : Datamarking, Delimiting, and Encoding. While these methods offered marginal mitigation, the average Attack Success Rates (ASR) remained significantly high. The Sandwich Defense yielded an average ASR of 0.82. Similarly, regarding the Spotlighting variants, we observed average ASRs of 0.52 for Datamarking, 0.74 for Delimiting, and 0.84 for Encoding. This failure suggests that structural formatting constraints cannot effectively override the backend LLM’s semantic understanding of malicious instructions.
Regarding detection, we assessed stealth capabilities through Perplexity (PPL) and Window PPL metrics (Liu et al., 2024). In contrast to real-world MCP tools (Fei et al., 2025), which show elevated perplexity levels of 1423.27 (PPL) and 1423.31 (Window PPL), QueryIPI maintains superior fluency with a PPL of 59.04 and a Window PPL of 130.98. By keeping perplexity far below the baseline of standard tool interactions, QueryIPI proves robust against statistical anomaly detection.
6 Conclusion
7 Limitations
As a general paradigm, query-agnostic attacks are not inherently restricted to the domain of coding agents. However, we limit our scope to coding agents in this work, given that our specific threat model operates under the assumption that the adversary possesses the capability to introduce custom tools into the victim agent’s toolset. Meanwhile, our method relies on specific architectural prerequisites: the target system must support user-defined tools and directly incorporate their descriptions into the agent’s context. Consequently, this attack is not limited to coding agents but theoretically extends to any framework satisfying these conditions. Operationally, the attack necessitates a capable backend LLM. While safety guardrails present a potential obstacle, they do not constitute a fundamental barrier. Adversaries can leverage established jailbreaking techniques (Wei et al., 2023; Zou et al., 2023) to bypass alignment filters, ensuring the practicality of the proposed threat model.
8 Ethical considerations
Appendix A Iterative Optimization Procedure
Algorithm 1 Iterative Optimization Procedure
2:Output: Optimal malicious description.
[…]
4:for $g=1$ to $G$ do
5:  for each description $d$ in $D_{\text{new}}$ do
6:   $s_{d}\leftarrow Scoring(d)$
7:   // Early stop mechanism
8:   if $s_{d}==FullScore$ then return $d$
9:   end if
10:   $Scores[d]\leftarrow s_{d}$
11:  end for
12:  // Choose the top score
13:  $(d_{\text{seed}},s_{\text{seed}})\leftarrow\text{SelectTop}(Scores)$
14:  $Scores\leftarrow\{\}$
15:  $Scores[d_{\text{seed}}]\leftarrow s_{\text{seed}}$
16:  $D_{\text{new}}\leftarrow\text{Mutation}(d_{\text{seed}},N)$
17:end for
18:
19:return $d_{\text{seed}}$
Appendix B Scoring Function
For each candidate tool description provided as input, the scoring function first injects it into the target agent. Subsequently, it systematically feeds all user queries from our training dataset to this modified agent and collects the corresponding responses for evaluation.
In designing our scoring mechanism, we deliberately moved away from a binary success/failure classification (e.g., 0 or 1) as seen in methodologies like PAIR (Chao et al., 2025). Instead, we adopted a more granular, continuous scoring approach, similar in spirit to the one used in TAP (Mehrotra et al., 2024), to capture nuanced outcomes.
In this task, a score of 0 indicates that the agent’s response for a given user query is completely irrelevant to the predetermined malicious action. Conversely, a score of 100 signifies a successful attack, where the response executes the predetermined malicious action. Intermediate scores represent states where the agent was influenced by the malicious tool description but the action was not fully completed, or where the agent’s defense mechanisms were triggered. To ensure the Judge LLM’s accuracy, we calibrated it using human feedback. Annotators ranked a set of generated agent outputs, and we then adjusted the prompt of Judge LLM to ensure its scoring rank order closely matched these human-provided rankings.
We performed a statistical analysis on 152 data points from our experiments to assess the alignment between the LLM judge and human evaluators. We observed an average Spearman’s $\rho$ of 0.7790 and an average Kendall’s $\tau$ of 0.6975. According to (Evans, 1996), these values indicate a strong correlation. Therefore, we conclude that the LLM judge demonstrates a high level of consistency with human judgment.
Appendix C Definition of $\mathcal{D}$
Hence, the overall search space can be written as
Appendix D RQ5: Cross-LLM Transferability
RQ5: How well does QueryIPI transfer from one backend LLM to other backend LLMs?
@@ Appendix D RQ5: Cross-LLM Transferability
To answer RQ5, we evaluated the cross-model generalization of our attack. We used the malicious tool description generated in RQ1 with 8 training samples (‘num=8’) and tested it on the five simulated coding agents. For a direct comparison, we also evaluated the TIP baseline under the same conditions. In this experiment, we systematically varied the agents’ backends, configuring them to use three different powerful, closed-source models: GPT-5, Grok-4, and Gemini-2.5-pro. The comparative results are presented in Table 4.
Our experimental results indicate that QueryIPI possesses a notable degree of transferability across different backend LLMs. As shown in Table 4, QueryIPI achieved an average Attack Success Rate (ASR) of 26% across all configurations, which is more than double the 10% average ASR of the TIP baseline. However, the ASR for both methods fluctuates depending on the agent-LLM combination. For instance, while QueryIPI achieved a remarkable 95% success rate on the Cline agent with a GPT-5 backend, the TIP baseline also saw its peak performance on this combination, albeit at a lower 55% ASR. Conversely, the attack on Copilot proved to be brittle and non-transferable, with its ASR plummeting from 85% to near 0% when the backend LLM was switched. We attribute the poor transferability to Copilot’s unstructured prompt format where all tool definitions are concatenated into a single line, necessitating model-specific tuning for the attack. Yet, this structural obfuscation is a double-edged sword: while it breaks fragile exploits upon model updates, it likely degrades standard performance by complicating tool parsing.
@@
Appendix E Prompt
E.1 Mutation Prompt